# Decoding Decoded: Understanding Hyperparameter Effects in Open-Ended Text Generation

### v1

@@ Abstract
Decoding strategies for large language models (LLMs) are a critical but often underexplored aspect of text generation tasks. Since LLMs produce probability distributions over the entire vocabulary, various decoding methods have been developed to transform these probabilities into coherent and fluent text, each with its own set of hyperparameters. In this study, we present a large-scale, comprehensive analysis of how hyperparameter selection affects text quality in open-ended text generation across multiple LLMs, datasets, and evaluation metrics. Through an extensive sensitivity analysis, we provide practical guidelines for hyperparameter tuning and demonstrate the substantial influence of these choices on text quality. Using three established datasets, spanning factual domains (e.g., news) and creative domains (e.g., fiction), we show that hyperparameter tuning significantly impacts generation quality, though its effects vary across models and tasks. We offer in-depth insights into these effects, supported by both human evaluations and a synthesis of widely-used automatic evaluation metrics.
@@
Figure 1: The quality of machine-generated text, as assessed by both automatic metrics and human judgment, is strongly influenced by hyperparameter selection.
@@ 1 Introduction
Large language models (LLM) do not readily output natural language text, but instead provide the user only with a high-dimensional probability distribution. The process of transforming these probabilities into text is referred to as decoding and can have nearly as much of an impact on the quality of the generated text as the LLM backbone itself. Several decoding strategies currently employed in conjunction with contemporary state-of-the-art (SOTA) LLMs encompass hyperparameters playing critical roles and hence significantly influencing the quality of the generated text. However, despite their importance, the selection and tuning of these hyperparameters remain an under-explored area, with the users often resorting to defaults and rather benchmarking different models instead of decoding strategies. In particular, this neglects differences in the desired properties of the generated texts, ranging from factual domains, like news generation, to more creative ones like (fiction) literature. Further, the success of decoding strategies might differ between models, a phenomenon that cannot be uncovered by the current practice of only assessing strategies globally, i.e. across models. Recent research emphasizes that different hyperparameter configurations can greatly impact both the coherence and diversity of generated text, underscoring the need for systematic evaluation Zhou et al. (2024).
In this study, we systematically vary hyperparameters of commonly used decoding strategies to assess their effect on text quality across different data sets and importantly, across different SOTA open-source models. By using pre-defined grids for the hyperparameters in deterministic (beam width), stochastic ($k$, $p$, and temperature), and contrastive ($\alpha$ and $k$) decoding strategies, we explore their role in balancing creativity, trustworthiness, and overall text quality. This is especially pertinent in addressing the common problem of degeneration and ensuring effective adaption to diverse text generation tasks Shi et al. (2024). Moreover, modern LLMs, while highly advanced, are still prone to issues like incoherence or degeneration under various hyperparameter settings. Our work addresses these challenges by providing a sensitivity analysis, offering practical guidelines for tuning decoding hyperparameters to optimize performance while mitigating degeneration
@@
This study advances the field of decoding strategies for LLMs by performing a large-scale sensitivity analysis of commonly used methods and examining their practical effects on model performance.
@@
    We conduct an extensive experimental analysis to evaluate the impact of hyperparameter values on various text quality metrics in open-ended text generation.
@@
    In total, we generate 2.2 million text continuations, which are publicly available for future meta-analyses, along with our complete codebase: https://github.com/YecanLee/Decoding-Decoded
@@ 2 Decoding Strategies and Hyperparameters
Decoding strategies are generally categorized into two types: deterministic and stochastic. Given the complexity introduced by their hyperparameters, we additionally separate contrastive strategies (cf. Sec. 2.3) from the other deterministic ones (cf. Sec. 2.1). While the latter is solely focused on maximizing the joint probability, the former explicitly compromise this objective by incentivizing more diverse texts.
@@ 2.1 Deterministic Strategies
This strategy selects the token with the highest probability at each time step, yielding a sequence that maximizes the local likelihood. However, it often leads to suboptimal results due to the lack of look-ahead, which can cause the model to get trapped in less coherent output.
@@
Freitag and Al-Onaizan (2017) extends the greedy approach by maintaining a beam of the $w$ most probable sequences at each time step. The hyperparameter $w$, known as the beam width, controls the trade-off between diversity and computation. Beam search reduces the risk of suboptimal choices but can still generate repetitive or generic text in open-ended tasks.
Frustratingly Simple Decoding (FSD).
Yang et al. (2024b) propose a method that contrasts outputs from the LLM with an auxiliary anti-LM, constructed using the current prefix. This method seeks to amplify the distinctions between probable and improbable tokens, yielding more robust generation in certain contexts.
2.2 Sampling Strategies
@@
Introduced by Ackley et al. (1985), this method samples from the full token distribution while modifying the softmax input. The temperature hyperparameter controls the sharpness of the distribution—higher temperatures flatten it, increasing randomness, while lower temperatures make it more deterministic, favoring higher-probability tokens.
@@
Proposed by Fan et al. (2018), this strategy limits sampling to the top $k$ tokens with the highest probabilities, forming a subset $V^{(k)}$. By truncating the distribution, it reduces the risk of sampling from low-probability (and potentially incoherent) tokens, while still introducing some diversity.
@@
Also known as $p$-sampling, this method was introduced by Holtzman et al. (2019). It samples from the smallest subset $S$ whose cumulative probability exceeds a threshold $p$. Nucleus sampling adapts dynamically to the token distribution, preserving both diversity and relevance by considering only the most probable tokens.
Typical sampling.
Meister et al. (2023) developed this method to produce more diverse and interesting text by aiming for tokens with an expected level of information content. It avoids both overly probable and highly improbable tokens, resulting in a more balanced generation suited for open-ended contexts.
@@ 2.3 Contrastive Strategies
Su et al. (2022) introduce a look-ahead mechanism that penalizes tokens disrupting the isotropy of the latent space in the language model, leading to more coherent and semantically consistent text.
Contrastive Decoding.
Li et al. (2023) propose searching for tokens that maximize the difference in probability between a large language model (LLM) and a weaker auxiliary model, helping to refine generation quality by selecting tokens that best distinguish high-quality text.
Adaptive Contrastive Decoding.
(Zhu et al., 2024) and Adaptive Contrastive Search. (Garces-Arias et al., 2024) propose adaptive strategies that adjust to model uncertainty (measured by Shannon entropy) at each time step. These methods aim to strike a balance between coherence and diversity dynamically, removing the need for extensive hyperparameter tuning.
@@ 3 Related Work
The choice of decoding strategy plays a crucial role in shaping the performance of text generation models, significantly affecting the trade-offs between output quality and diversity. Wiher et al. (2022) provide a foundational analysis of decoding methods such as beam search, top-k sampling, and nucleus sampling. Their findings indicate that while beam search is effective for structured tasks like machine translation, it often leads to repetitive or less coherent text in creative tasks like story generation. This highlights the need to adapt decoding strategies to specific task requirements.
Subsequent research, including Amini et al. (2023), builds on these insights by offering a comprehensive review of the principles guiding text generation. The study emphasizes how various decoding mechanisms influence the final output, underscoring the necessity of selecting appropriate strategies based on the task at hand. It also explores the balance between fluency and diversity, particularly in open-ended text generation tasks.
More recently, Shi et al. (2024), investigates how decoding strategies scale with large language models (LLMs). This work explores the interaction between decoding strategies, hyperparameters such as top-k and temperature, and model size. The authors illustrate the increasing complexity of optimizing these strategies as models grow in scale, particularly in open-ended generation scenarios.
Additionally, Zhou et al. (2024) offers an in-depth exploration of sampling-based methods. It provides guidelines on managing the balance between diversity and the risk of incoherence, demonstrating how hyperparameter tuning can shape the quality and diversity of generated outputs. This work serves as a practical resource for selecting decoding strategies based on task-specific needs.
However, most existing studies fail to provide an extensive and detailed analysis of how various quality metrics, such as coherence, diversity, and MAUVE, are affected by different decoding strategies. To address this gap, we introduce a new metric, QText, which combines these metrics using the harmonic mean. We compare the best and worst performing strategies based on QText with those identified by human evaluators, highlighting areas of agreement and divergence. We provide detailed insights in Section A2.
@@ 4 Experimental setup
We employ seven models to generate stories from prompts sourced from three distinct datasets, using six decoding strategies with varying hyperparameter configurations. To evaluate the quality of the generated text, we rely on three widely used automatic metrics: coherence, diversity, and MAUVE. In addition, human evaluators are engaged to provide further assessment of the quality of the generations.
@@ 4.3 Evaluation Metrics
In this study, we base our metric evaluations on multiple datasets, yielding a total of 1,242 measurements.
Human Eval.
In order to evaluate the quality of the generated text, we consider two critical aspects: fluency, coherence and similarity to human text. A fluent piece of text is written in grammatical English and has a natural flow (e.g. excluding unnatural repetition or web formatting). A coherent piece of text should stay on topic with the prompt and avoid unnatural topic drift. We provide five native English speakers with 39 competing continuations for three prompts and ask them to rate them based on their perceived quality. Definitions and instructions for the rating process are shown in Appendix A4, Figure 27.
@@ 4.4 Datasets
We evaluate our proposed method across three domains in open-ended text generation: news, Wikipedia, and stories. For the news domain, we utilize 2,000 articles from Wikinews; for the Wikipedia domain, we draw from the WikiText-103 dataset (1,314 examples; Merity et al., 2016); and for the stories domain, we employ the Project Gutenberg split of BookCorpus (1,947 examples; Zhu et al., 2015). Each example consists of a prompt paired with a gold-standard, human-generated continuation for evaluation purposes.
For each prompt, we generate 256 tokens as a continuation. The resulting text is assessed using both automatic metrics (outlined in Section 4.3 ) and human evaluations.
@@ 5 Results
Table 2: Automatic evaluation results for all seven models across three datasets. Scores are aggregated as weighted averages, accounting for the different number of generations per dataset. The highest scores for each strategy are highlighted in bold, with overall best results additionally marked by a $\star$.
In this section, we present our main results. Table 2 shows combinations of models, decoding strategies, and hyperparameter selections, along with their automatic evaluations in terms of Diversity, MAUVE, and Coherence. The values are computed as weighted means across the three datasets.
The results confirm the sensitivity of text generation quality to the specific values chosen for decoding strategies. The highest-ranking strategies consistently achieve a balance between coherence and diversity, demonstrating that successful text generation often hinges on not favoring one quality at the expense of the other. In contrast, lower-ranking strategies tend to overemphasize one aspect—either coherence or diversity—leading to less optimal outputs. For example, beam search typically favors coherence at the cost of diversity, while contrastive search with high values of $\alpha$ and $k$ prioritizes diversity, resulting in less coherent text.
Interestingly, the results also reveal that smaller models, such as GPT2-XL (1.5B), can outperform larger models like Falcon2 (11B) when the correct decoding strategy is applied. This finding underscores the importance of selecting optimal hyperparameters and decoding methods, which can significantly impact performance, regardless of model size.
In terms of human evaluation, there is a moderate correlation with automatic evaluation metrics, but the variability suggests that existing automatic metrics may not fully capture the nuanced preferences of human raters. As previously noted by Su and Collier (2023) and Garces-Arias et al. (2024), metrics like MAUVE tend to favor diverse text outputs, even when coherence decreases. Our findings confirm this observation, as human judgments often align more closely with a balance between coherence and diversity, rather than with the extremes favored by individual metrics. Nevertheless, as shown in Figures 2 and 3, there are similarities in the extremes of the spectrum: Both automatic scoring through QText and human scoring show preferences for decoding strategies that balance coherence and diversity better, such as contrastive search with moderate values of $\alpha$ and $k$, adaptive contrastive search, and sampling methods that allow diversity. A comprehensive overview of hyperparameter areas of high performance is presented in Table 2 and more detailed in Figures 4, 6, 19, 20, 22 and 24.
Finally, the moderate correlation between human evaluations and automatic metrics, as presented in Figure 26, underscores the need for more reliable and robust automatic evaluation methods. While current metrics offer valuable insights, they fall short of accurately reflecting the quality of open-ended text generation as perceived by human evaluators. This highlights the ongoing challenge of developing automatic metrics that can more effectively balance the competing factors of coherence and diversity in open-ended text generation tasks.
@@
Figure 2: Top five and bottom five decoding strategies across datasets, based on human scores. The highest-ranking strategies generally strike a balance between coherence and diversity, while the lowest-ranking strategies tend to overemphasize one at the expense of the other—such as beam search, which favors coherence, or contrastive search with $\alpha=1.0$ and $k=50$, which prioritizes diversity.
Figure 3: Top five and bottom five decoding strategies across datasets, based on QText. The score exhibits similarities to human preferences in Figure 2.

### v2

@@ Abstract
Decoding strategies for generative large language models (LLMs) are a critical but often underexplored aspect of text generation tasks. Guided by specific hyperparameters, these strategies aim to transform the raw probability distributions produced by language models into coherent, fluent text. In this study, we undertake a large-scale empirical assessment of a range of decoding methods, open-source LLMs, textual domains, and evaluation protocols to determine how hyperparameter choices shape the outputs. Our experiments include both factual (e.g., news) and creative (e.g., fiction) domains, and incorporate a broad suite of automatic evaluation metrics alongside human judgments. Through extensive sensitivity analyses, we distill practical recommendations for selecting and tuning hyperparameters, noting that optimal configurations vary across models and tasks. By synthesizing these insights, this study provides actionable guidance for refining decoding strategies, enabling researchers and practitioners to achieve higher-quality, more reliable, and context-appropriate text generation outcomes.
@@
Figure 1: Influence of the nucleus sampling hyperparameter $p$ on the distribution of diversity and coherence metrics in text generated by Mistral 7B v0.3 (green). For comparison, the distribution of the same metrics in human-written text is displayed in blue.
@@ 1 Introduction
Generative large language models (LLMs) do not directly produce natural language text. Instead, they generate a high-dimensional probability distribution over all tokens in their vocabulary. The process of converting these probabilities into coherent text, known as decoding, can substantially influence the quality of the generated output, sometimes matching the impact of the LLM itself.
Most decoding strategies employed with contemporary LLMs involve hyperparameters that play critical roles in shaping the generated text. These hyperparameters strongly influence factors such as coherence, fluency, and diversity (Zhou et al., 2024). Despite their importance, the selection and tuning of these hyperparameters remain under-explored. Users often rely on default settings and prioritize benchmarking different models over optimizing decoding strategies. This approach overlooks the varying requirements of different text generation tasks, ranging from factual domains like news generation to creative areas such as fiction. Additionally, the effectiveness of decoding strategies may vary across different models, a nuance that current practices— which assess strategies uniformly across models—fail to capture.
Recent research underscores the impact of hyperparameter configurations on both the coherence and diversity of generated text when using sampling-based decoding (Zhou et al., 2024). In this study, we extend this by systematically varying the hyperparameters of commonly used decoding strategies—deterministic, sampling-based, and contrastive—to evaluate their effects on text quality across diverse datasets, metrics, and open-source LLMs. By employing predefined grids for hyperparameters in deterministic (beam width), stochastic (top-$k$, top-$p$, and temperature), and contrastive ($\alpha$ and $k$) decoding strategies, we investigate their role in balancing coherence, fluency, diversity, and overall text quality. This investigation is particularly relevant for addressing common issues such as degeneration and ensuring effective adaptation to a wide range of text generation tasks Shi et al. (2024).
Furthermore, while modern LLMs are highly advanced, they remain susceptible to problems like incoherence or degeneration under certain hyperparameter settings. Our work addresses these challenges by providing a sensitivity analysis and offering practical guidelines for tuning decoding hyperparameters. This aims to optimize the desired properties of the generated text while mitigating undesirable behaviors, including repetitiveness, and inconsistencies or hallucinations.
@@
This study advances the field of decoding strategies for LLMs by performing a large-scale sensitivity analysis of commonly used decoding methods and examining their practical effects on model performance.
@@
    We conduct an extensive experimental analysis to evaluate the impact of hyperparameter values on various text quality metrics across models in open-ended text generation.
@@
    We create (and share) a unique database for future research: In total, we generate 2.2 million text continuations, which are publicly available for future meta-analyses, along with our complete codebase: https://github.com/YecanLee/Decoding-Decoded.
@@ 2 Decoding Strategies and Hyperparameters
Decoding strategies are generally categorized into two types: deterministic and stochastic. Given the complexity introduced by their hyperparameters, we additionally separate contrastive strategies (cf. Sec. 2.3) from the other deterministic ones (cf. Sec. 2.1). While the latter are solely focused on maximizing the joint probability, the former explicitly compromise this objective by incentivizing more diverse texts. We further do not go into detail on other existing strategies that we do not employ, such as Frustratingly Simple Decoding (Yang et al., 2024b), Typical sampling (Meister et al., 2023), Contrastive Decoding (Li et al., 2023), and Adaptive Decoding (Zhu et al., 2024).
@@ 2.1 Deterministic Strategies
This strategy selects the token with the highest probability at each time step, resulting in a sequence that maximizes the local likelihood. However, it often leads to suboptimal results due to the lack of look-ahead, which can cause the model to get trapped in less coherent output.
@@
Freitag and Al-Onaizan (2017) extends the greedy approach by maintaining a beam of the $w$ most probable sequences at each time step. The hyperparameter $w$, known as the beam width, aims to reduce the risk of suboptimal sequence choices, but the results can still generate repetitive text in open-ended tasks.
2.2 Sampling-based Strategies
@@
Introduced by Ackley et al. (1985), this method samples from the full distribution over all tokens while modifying the softmax. The temperature hyperparameter controls the sharpness of the distribution — higher temperatures flatten it, increasing randomness, while lower temperatures make it more deterministic, favoring tokens with higher probabilities even more.
@@
Proposed by Fan et al. (2018), this strategy limits sampling to the top $k$ tokens with the highest probabilities, forming a subset $V^{(k)}$. Truncating the distribution reduces the risk of sampling from the long tail of low-probability (and potentially incoherent) tokens, while still introducing some diversity through stochasticity.
@@
Also known as top-$p$ sampling, this method was introduced by Holtzman et al. (2019). It samples from the smallest subset $S$ whose cumulative probability exceeds a threshold $p$. Nucleus sampling adapts dynamically to the token distribution, preserving both diversity and relevance by considering only the most probable tokens.
@@ 2.3 Contrastive Strategies
Su et al. (2022) introduce a look-ahead mechanism that penalizes tokens disrupting the isotropy of the latent space in the language model, penalizing degeneration while producing more semantically consistent text.
Adaptive Contrastive Search.
Garces-Arias et al. (2024) propose an adaptive strategy that aims to strike a balance between coherence and diversity dynamically, based on the model entropy at each time step, removing the need for extensive hyperparameter tuning.
@@ 3 Related Work
The selection of decoding strategies is critical in determining the performance of text generation models, particularly in balancing output quality and diversity. Wiher et al. (2022) provide a foundational analysis of decoding methods, including beam search, top-$k$ sampling, and nucleus sampling. Their findings indicate that while beam search is effective for structured tasks such as machine translation, it often results in repetitive or less coherent text in creative tasks like story generation. This underscores the necessity of adapting decoding strategies to specific task requirements.
Subsequent studies, such as Amini et al. (2023), build on these insights by offering a comprehensive review of the principles that guide text generation. This research highlights the influence of various decoding mechanisms on the final output, emphasizing the importance of selecting suitable strategies tailored to specific tasks. Additionally, it examines the trade-off between fluency and diversity, particularly in open-ended text generation scenarios. More recently, Shi et al. (2024) investigate how decoding strategies scale with large language models (LLMs), exploring the interactions between decoding strategies, their hyperparameters, and model size. The authors demonstrate the increasing complexity of optimizing these strategies as models grow in scale, especially in open-ended generation contexts. Our experiments substantially extend the work of Shi et al. (2024) by exploring additional hyperparameter combinations across a broader range of models and evaluating text quality using multiple metrics beyond MAUVE.
Additionally, Zhou et al. (2024) provide an in-depth exploration of sampling-based methods. They offer guidelines for managing the balance between diversity and the risk of incoherence, illustrating how hyperparameter tuning can influence the quality and diversity of generated outputs. This work is a practical resource for selecting decoding strategies based on task-specific requirements. In contrast, our focus extends beyond the evaluation of risk associated with tuning temperature and truncation parameters in stochastic strategies.
While providing valuable insights, many existing studies do not offer a comprehensive analysis of how different decoding strategies affect various quality metrics, such as coherence, diversity, and MAUVE. We address this gap and introduce QText, which combines these metrics using the harmonic mean. We compare the best and worst-performing strategies based on these metrics with those identified by human evaluators, highlighting areas of agreement and divergence. Detailed insights are provided in Figure 3 and Section A.2.
@@ 4 Experimental setup
We employ seven models to generate stories from prompts sourced from three distinct datasets, using six decoding strategies with varying hyperparameter configurations. To evaluate the quality of the generated text, we rely on three widely used automatic metrics: coherence, diversity, and MAUVE, and we also measure QText. In addition, human evaluators are engaged to provide further assessment of the quality of the generations.
@@ 4.3 Evaluation Metrics
Human Evaluation.
To evaluate the quality of the generated text, we consider two critical aspects: fluency and coherence. A fluent text is written in grammatical English and has a natural flow (e.g. excluding unnatural repetition or web formatting). A coherent text should stay on topic with the prompt and avoid unnatural topic drift. We provide five native English speakers with 39 competing continuations for one prompt per dataset, and subsequently ask them to rank them based on their perceived quality, for a total of 570 evaluations. Definitions and instructions for the rating process are shown in Appendix A.4, Figure 26.
@@ 4.4 Datasets
We evaluate our proposed method across three domains in open-ended text generation: news, Wikipedia, and stories. For the news domain, we utilize 2,000 articles from Wikinews; for the Wikipedia domain, we employ 1,314 examples from the WikiText-103 dataset (Merity et al., 2016); and for the stories domain, we employ the Project Gutenberg split of BookCorpus (1,947 examples; Zhu et al., 2015). Each example consists of a prompt paired with a gold-standard, human-generated continuation for evaluation purposes. For each prompt, we generate 256 tokens as a continuation. The resulting text is assessed using both automatic metrics (outlined in Section 4.3 ) and human scores.
@@ 5 Results
Table 2: Aggregated automatic evaluation results: Average scores across all seven models and all three datasets. Weighted averages are used, accounting for the different sample sizes in the datasets. The highest scores per strategy are highlighted in bold, with the best results overall additionally marked by a $\star$.
@@
Figure 2: Top five and bottom five decoding strategies, based on QText averages for each dataset. The highest-ranking strategies generally strike a balance between coherence and diversity, while the lowest-ranking strategies tend to overemphasize one at the expense of the other—such as beam search, which favors coherence, or contrastive search with $\alpha=1.0$ and $k=50$, which prioritizes diversity.
Automatic Evaluation
Below, we present a comprehensive interpretation of the results reported in Table 2. These findings aggregate the performance of various decoding strategies and their associated hyperparameters across multiple models and datasets, taking into account weighted averages to reflect dataset size differences. The reported metrics include coherence, diversity, MAUVE, and QText. Notably, coherence has a theoretical maximum of 0, while diversity, MAUVE, and QText each have a theoretical maximum of 100. The final row represents human-produced text, with reference values of approximately -2.74 for coherence, 93.28 for diversity, 100 for MAUVE, and 87.37 for QText, serving as an aspirational benchmark for automated methods.
Beam search yields coherence values around -0.55 to -0.72, which are far from the human coherence of -2.74. Its diversity remains low, peaking at only 12.78 and dropping to 6.17 as the beam width increases. MAUVE and QText scores under beam search never exceed 25.27 and 21.50, respectively, well below the human benchmarks.
Contrastive search achieves its best performance with moderate values of $k$ and $\alpha$, such as $(\alpha=0.8,k=3)$ or $(\alpha=0.6,k=10)$. However, combinations involving higher values of $k$ and $\alpha$ tend to reduce coherence to levels considerably below human-like values, sometimes reaching as low as -5.76. In contrast, smaller values of the degeneration penalty, such as $\alpha=0.2$, are associated with very high coherence but extremely low diversity, resembling the performance characteristics of beam search. Even with optimal hyperparameter settings, achieving coherence and diversity levels comparable to human performance remains a challenge.
Adaptive contrastive search (hyperparameter-free) achieves strong overall performance. It reaches a diversity of 93.94 and a QText of 85.72, both near human levels. Its coherence is -1.68, which is closer to human than many contrastive search configurations, but still not ideal. MAUVE is 79.61, which is better than basic beam search but not as high as the best stochastic approaches.
Sampling with temperature exhibits a trade-off between coherence and diversity that becomes more favorable at higher temperatures. At a temperature of 1.0, for instance, diversity reaches 93.27, MAUVE is 91.17, and QText is 83.30, all nearing human references. Coherence is -3.07, which is closer to human coherence than beam search or low-temperature sampling.
Top-$k$ sampling improves diversity and MAUVE as $k$ grows. At $k=10$, diversity is 77.10, MAUVE is 85.80, and QText is 80.28, all improved over lower $k$ values. However, coherence remains around -1.90, which is not close enough to the human target.
Top-$p$ (nucleus) sampling offers strong performance when $p=0.95$. At this value, diversity is 87.60, MAUVE is 90.05, and QText is 85.31, all close to human scores. Coherence at this setting is -2.21, which is closer to the human reference than many other strategies. Figure 1 illustrates this effect over the distribution of coherence and diversity for Mistralv03 outputs.
No single decoding strategy achieves human-level scores across all metrics; however, certain configurations exhibit performance closely aligned with human references. Notable examples include top-$p$ sampling at $p=0.95$, adaptive contrastive search, and sampling with temperature at $t=1.0$, which emerge as particularly effective methods.
These findings reflect aggregated performance across three datasets and seven models. Detailed results for individual models and configurations are presented in Figures 3 through 24 and discussed in Section 6. Additionally, Figure 2 highlights the top five best- and worst-performing combinations of models and decoding strategies based on QText. A notable observation is the frequent appearance of older and smaller architectures, such as GPT2-XL (1.5B) with configurations like CS_(0.6, 5) or CS_(0.8, 3), among the top performers. This underscores the importance of hyperparameter selection over model size. Conversely, the lowest-performing configurations consistently overemphasized a single metric, such as beam search prioritizing coherence or contrastive search with high $\alpha$ values emphasizing diversity. The inclusion of much larger architectures, such as Falcon2 (11B), among the lowest-performing configurations further indicates that model size alone does not guarantee superior performance. Case studies demonstrating the impact of hyperparameter choices on text generation are provided in Tables 4, 5, and 6.
Dataset Model Decoding Strategy Human Score QText
Top five decoding methods per dataset
book Mistralv03 CS_(0.4, 10) 97.50 78.31
GPT2-XL ACS 97.50 86.03
Mistralv03 CS_(0.4, 20) 95.00 82.30
Mistralv03 TopP_0.95 95.00 85.97
Mistralv03 TopP_0.95 95.00 85.97
wikinews Mistralv03 CS_(0.6, 10) 97.50 87.48
Mistralv03 CS_(0.4, 10) 95.00 88.12
Mistralv03 TopK_50 95.00 87.10
GPT2-XL ACS 92.50 86.31
Mistralv03 TopP_0.95 92.50 87.19
wikitext GPT2-XL ACS 97.50 84.81
Mistralv03 CS_(0.6, 5) 97.50 86.84
Mistralv03 TopP_0.95 97.50 83.22
Mistralv03 TopK_50 95.00 84.89
Mistralv03 CS_(0.4, 20) 95.00 81.13
Bottom five decoding methods per dataset
book Mistralv03 CS_(0.8, 50) 2.50 5.13
Mistralv03 CS_(1.0, 50) 2.50 3.10
Mistralv03 CS_(0.2, 50) 5.00 44.27
Mistralv03 CS_(0.8, 10) 5.00 9.45
GPT2-XL CS_(0.8, 50) 5.00 5.13
wikinews Mistralv03 BS_20 2.50 26.08
Mistralv03 BS_50 5.00 17.41
Mistralv03 CS_(0.8, 50) 5.00 6.73
Mistralv03 CS_(1.0, 5) 5.00 8.99
Mistralv03 CS_(0.8, 50) 5.00 6.73
wikitext Mistralv03 CS_(1.0, 50) 2.50 9.80
Mistralv03 CS_(1.0, 5) 5.00 29.25
Mistralv03 CS_(0.8, 10) 5.00 38.43
Mistralv03 CS_(1.0, 50) 5.00 9.80
Mistralv03 CS_(0.8, 50) 5.00 13.19
Table 3: Top and bottom five models and decoding strategies for human evaluators. Results indicate a moderate, statistically significant positive correlation based on 570 evaluations, $r_{\text{Pearson}}=0.64$ (p-value < 2.2e-16).
Human Evaluation
Table 3 summarizes the top five and bottom five decoding strategies as ranked by human evaluation scores. The findings reveal a strong alignment between humans and QText for high-performing strategies across datasets.
Specifically, in the book dataset, strategies such as ACS and CS_(0.4, 10) achieve high human scores (97.5), which correspond closely to their QText values (86.03 and 78.31, respectively). Similarly, in the wikinews and wikitext datasets, strategies like CS_(0.6, 10) and CS_(0.6, 5) achieve high human ratings (97.5) and strong QText values (87.48 and 86.84, respectively). For lower-performing strategies, discrepancies between humans and QText are observed. In the book dataset, the strategy CS_(0.2, 50) receives a low human score (5.0) but a comparatively higher QText value (44.27). Nevertheless, both human evaluations and QText consistently indicate that strategies overly prioritizing diversity, such as CS_(1.0, 50), or coherence, such as BS_50, tend to produce low-quality text generations.
To examine the broader alignment between automatic metrics with human preferences, we analyze their correlation. Figure 25 illustrates that human scores have moderate, statistically significant positive correlations with QText and MAUVE. Coherence demonstrates a lower but statistically significant positive correlation, while diversity exhibits a very weak and non-significant correlation. These findings highlight limitations of automatic metrics, consistent with previous work by Su and Collier (2023) and Garces-Arias et al. (2024).